OpenStudy (funnyguy2):

The coordinates below represent two linear equations. How many solutions does this system of equations have? Line 1 x y –2 7 0 6 Line 2 x y 3 0 0 –3 A. 0 B. exactly 1 C. exactly 2 D. infinitely many

OpenStudy (funnyguy2):

what do i do after i find the slopes?

OpenStudy (funnyguy2):

6 - 7 / 0 - -2 -1 / 2 -0.5 or -1/2 -3 - 0 / 0 - 3 -3 / -3 1

OpenStudy (funnyguy2):

@acxbox22

OpenStudy (funnyguy2):

i can eliminate d

OpenStudy (acxbox22):

first find the equations of both lines

OpenStudy (funnyguy2):

what do you mean?

OpenStudy (acxbox22):

like you are given two points from each line now find its equation y=mx+b m=slope b=y intercept

OpenStudy (funnyguy2):

y=-1/2*x+6

OpenStudy (funnyguy2):

@acxbox22

OpenStudy (acxbox22):

yes correct @funnyguy2 now do the same for the second line

OpenStudy (funnyguy2):

okay

OpenStudy (funnyguy2):

y=-1*x+3

OpenStudy (acxbox22):

u found the slope for the second line above which is 1 and the y intercept is -3 becasue it intercepts the y axis at (0,-3) so we get y=x-3 not y=-x+3

OpenStudy (funnyguy2):

so how do i figure out if it's b or c?

OpenStudy (acxbox22):